A 41.5-N puck is traveling at 4.7m/s. It strikes a 11.6-N puck, which is stationary. The two pucks stick together. Give the common final speed of the two.


1
Expert's answer
2021-12-02T10:05:44-0500

"F=mg\to m_1=41.5/9.8=4.2\ (kg)"


"m_2=11.6/9.8=2.76\ (kg)"


"m_1v_1=(m_1+m_2)v_2\to v_2=m_1v_1/(m_1+m_2)="


"=4.2\cdot4.7/(4.2+2.76)=2.84\ (m/s)" . Answer